OpenStudy (jjuden):

Solve using the perfect squares method (2x+5)^2=25

hero (hero):

Square Root Both sides. For example, if you had \((x + 3)^2 = 9\) You would recognize that the left side is a binomial square and the right side contains an integer square. Then you would express the equation as: \((x + 3)^2 = 3^2\) If you recall \(3^2 = 3(3) = 9\) Then you would take the square root of both sides: \(\sqrt{(x + 3)^2} = \sqrt{3^2}\) Square root and square are inverses of each other. In other words, they cancel. So: \(x + 3 = 3\) And then you isolate x afterwards: x = 3 - 3 x = 0

hero (hero):

Also keep in mind that for your problem, there are two answers since \(\sqrt{a^2} = \pm a\)